# Event-Based Tracking Any Point with Motion-Augmented Temporal Consistency

### v1

@@ Abstract
Tracking Any Point (TAP) plays a crucial role in motion analysis. Video-based approaches rely on iterative local matching for tracking, but they assume linear motion during the blind time between frames, which leads to target point loss under large displacements or nonlinear motion. The high temporal resolution and motion blur-free characteristics of event cameras provide continuous, fine-grained motion information, capturing subtle variations with microsecond precision. This paper presents an event-based framework for tracking any point, which tackles the challenges posed by spatial sparsity and motion sensitivity in events through two tailored modules. Specifically, to resolve ambiguities caused by event sparsity, a motion-guidance module incorporates kinematic features into the local matching process. Additionally, a variable motion aware module is integrated to ensure temporally consistent responses that are insensitive to varying velocities, thereby enhancing matching precision. To validate the effectiveness of the approach, an event dataset for tracking any point is constructed by simulation, and is applied in experiments together with two real-world datasets. The experimental results show that the proposed method outperforms existing SOTA methods. Moreover, it achieves 150% faster processing with competitive model parameters. The project page is here.
@@ 1 Introduction
Figure 2: (a) The spatio-temporal distribution of events generated by a stick rotating uniformly around a pivot. Sampling from different patches along the stick reveals spatial sparsity of events. Counting the events at each patch shows a positive correlation between event update frequency and object speed, which causes temporal inconsistencies with varying motion speeds. (b) This paper leverages the temporal continuity of events to capture subtle motion changes, enhancing the matching process and thereby improving tracking accuracy.
Recent methods rely on video input, predicting the positions of query points by matching their appearance features with local regions in subsequent frames [9, 10, 15]. However, as these methods assume slow linear motion during the blind time between frames, they faces the challenge of objects may undergo large displacements or nonlinear motion, causing query points to exceed the bounds of local regions and resulting in ambiguities in feature matching, see Fig. 1. While some approaches attempt to mitigate this by considering spatial context [36, 7], they still struggle to overcome the lack of motion during blind time.
To cope with the above issue, this paper utilize event cameras to capture the motion during blind time. Event cameras [22, 33] are bio-inspired sensors that respond to pixel-level brightness changes with microsecond temporal resolution, generating sparse and asynchronous event streams. They have the characteristics of high temporal resolution, no motion blur, and low energy consumption. Therefore, parsing motion during blind times using event streams is a feasible solution for tracking any point. Furthermore, capturing motion alone significantly reduces computational overhead compared to traditional frame-based cameras, enabling more efficient methods.
However, as shown in Fig. 2(a), the unique spatio-temporal properties of events make existing methods difficult to apply, manifesting in two ways: 1) Event cameras respond only to pixels with brightness changes, resulting in a sparse spatial distribution that introduces ambiguities in appearance-based matching due to spatial discontinuities. 2) The variable speed of scene objects causes fluctuations in event update frequencies, impacting the temporal consistency of event representations. Higher motion speeds lead to increased event update frequencies and densities, and vice versa. Density variations cause inconsistencies in event representations over time, affecting the precision and reliability of temporal matching.
@@
Therefore, this paper proposes a novel event-based point tracking framework that comprises a Motion-Guidance Module (MGM) and a Variable Motion Aware Module (VMAM). Specifically, MGM leverages the gradient from event time surface to compute kinematic features that construct a dynamic-appearance space, clarifying ambiguities in feature matching and guiding the extraction of temporally consistent appearance features. Additionally, to model the non-stationary states of events, VMAM is introduced to employ kinematic cues for adaptive correction and generate temporally consistent responses by combining long-term memory parameters with hierarchical appearance. The method is evaluated on a synthetic dataset as well as two real-world datasets, with experimental results demonstrating its superiority. The main contributions of this paper can be summarized as follows:
  • An event-based framework for tracking any point is presented to monitor surface points on objects by leveraging the temporal continuity of events.
  • This paper reveals the impact of spatial sparsity and motion sensitivity in event data on TAP. To address these limitations, a motion-guidance module is proposed to enhance matching process using temporal continuity, while a variable motion aware module models kinematic cues to estimate non-stationary event states, thereby improving tracking accuracy.
  • Experimental results demonstrate that the proposed approach outperforms current state-of-the-art methods on both synthetic and real-world datasets. Moreover, the proposed method exhibits competitive model parameters and 150% faster computational speed.
@@ 2.1 Event-based Optical Flow
Figure 3: (a) Framework overview. Given the event data and the initial positions of target points as input, the model initializes the locations for subsequent time steps, along with appearance features. It then iteratively calculates kinematic features and updates the appearance correlation map at each point to refine the trajectory. (b) Motion-Guidance Module. MGM extracts kinematic features from the gradient information in the event stream, guiding appearance feature matching and forming a dynamic-appearance matching space with the appearance features. (c) Variable Motion Aware Module. VMAM leverages kinematic features from MGM to produce temporally consistent feature responses, thereby resulting in robust correlation maps.
@@ 2.2 Event-based Feature Tracking
Feature tracking aims to predict the trajectories of keypoints. Early approaches can be grouped into two types: one [20, 39] treats feature points as event sets and tracks them using the ICP [5] method, while the other [11] extracts feature blocks from reference frames and matches them by computing brightness increments from events. Moreover, event-by-event trackers [2, 3, 18] exploit the inherent asynchronicity of event streams. Unfortunately, these methods involve complex model parameters that require extensive manual tuning for different event cameras and new environments.
@@ 2.3 Tracking Any Point
Tracking any point based on events has yet to be proposed, while techniques for using standard frames have been developed [9, 10, 15, 7, 38]. These methods model the appearance around the points, using MLPs to capture long-range temporal contextual relationships across frames. During inference, they employ a sliding time window to handle long videos. For example, PIPs [15] frames pixel tracking as a long-range motion estimation problem, updating trajectories through iterative local searches. In contrast, TAP-Net [9] formalizes the problem as tracking any point, overcoming occlusion through global search. However, these methods track points independently, leading to ambiguities in feature matching. Consequently, some studies [36, 7, 19] have been proposed to utilize spatial context to alleviate this issue. In addition to methodological innovations, the PointOdyssey dataset [38] is collected to advance the field, featuring the longest average video length and the highest number of tracked points to date.
Unfortunately, applying these methods to event data encounters several limitations. The spatial sparsity of events leads to misalignment when solely modeling based on appearance. Additionally, the temporal inconsistency in event density, caused by variable-speed motion, affects temporal context modeling. In this paper, a motion-guidance module is designed to construct a dynamic-appearance matching space, thereby reducing matching ambiguity. Moreover, a variable motion aware module is employed to generate temporally consistent responses for correlation operations. The method is trained and tested on simulated event modalities derived from the [38] dataset.
@@ 3.1 Setup and Overview
Specifically, let $E_{j}=\left\{\left(x_{k},y_{k},t_{k},p_{k}\right)\right\}_{k=1}^{N}$ denote the event stream from $t_{j-1}$ to $t_{j}$, where $N$ is the number of events, and each event is a 4-tuple consisting of the coordinates $x_{k}$ and $y_{k}$, timestamp $t_{k}$, and polarity $p_{k}\in\{-1,+1\}$. Given a target point $x_{src}\in\mathbb{R}^{2}$, subsequent events are represented by the Time Surface (TS) [26], where each pixel records the timestamp of the most recent event, capturing motion process over a period of time.. During the initialization period, the TS representation is fed into a residual network [16] to extract features $\left\{F_{0},F_{1},...,F_{t}\right\}$. The point trajectories at all time steps are initialized as:
$\centering X^{0}=\left\{x_{0}^{0},x_{1}^{0},...,x_{t}^{0}\right\}=\left\{x_{src},x_{src},...,x_{src}\right\},\@add@centering$ (1)
with the corresponding feature at time $t$ being $f_{t}^{0}=F_{t}(x_{t}^{0})$.
At the iterative stage, let $x_{t}^{k}$ represent the coordinate of the query point at time $t$ after the $k$-th iteration. This paper employs VMAM to model the non-stationary states of point features at times $t-2$, $t-4$, and the initial moment, leveraging these to compute correlations $c_{t}^{k}$ with the neighboring features around $x_{t}^{k}$ at time $t$. A transformer takes correlations $C^{k}$, the kinematic features $V^{k}$ derived from MGM, along with the position-encoded apparent point motions $X_{t}^{k}-X_{t-1}^{k}$ as input to obtain the point displacement $\Delta X$. Subsequently, the point coordinates are updated through Eq. 2:
$\centering X^{k+1}=X^{k}+\Delta X,\@add@centering$ (2)
and the process is repeated iteratively.
@@ 3.2 Motion-Guidance Module
To reduce ambiguities arising from appearance-only matching, a motion-guidance module is designed to leverage the gradient characteristics of the TS to compute kinematic features, thereby building a dynamic-appearance matching space and subsequently guiding the extraction of temporally consistent appearance features.
Specifically, the event stream after TS encoding is visualized as a surface in the $xyt$ spacetime domain, representing the active events surface $\Sigma_{e}$ [4]. The spatial gradients of this surface describe the temporal changes relative to spatial variations, establishing a derivative relationship with pixel displacement at corresponding positions, see Eq. 3
@@
For the query point $x_{t}^{k}$, this paper treats the surface formed by neighboring pixels as a surface of active events. Using the coordinates of neighboring points $(x_{1},y_{1},t_{1})$, $(x_{2},y_{2},t_{2})$, …, a system in Eq. 4 is constructed:
@@
Here, $(a,b,c,d)$ are coefficients for the tangent plane. The spatial gradient of the surface is estimated through plane fitting using SVD, providing kinematic vectors at the target pixel. However, in the presence of overlapping moving objects, the pixels at boundary intersections can exhibit multiple distinct motion states, disrupting local smoothness and resulting in deviations in the kinematic vectors. To address this, the proposed motion-guidance module employs a multi-layer perceptron to capture the temporal motion relationships, dynamically assigning weights to the kinematic features at different time steps, thereby correcting the motion ambiguity at the boundaries.
The corrected kinematic features serve two key roles: first, as input to VMAM, guiding the generation of speed-insensitive appearance features for correlation; second, as inputs to the transformer, combining with correlation maps and position-encoded point motions to construct a dynamic-appearance space for iterative point displacement updates.
@@ 3.3 Variable Motion Aware Module
Objects in the scene exhibit variable speeds in two primary ways: either different objects move at distinct speeds or individual objects vary their speed over time. The variation in speed affects the frequency of event updates. Assuming an object moves at speed $\boldsymbol{v}=(v_{x},v_{y})$, the illuminance change rate at any point on the edge is given by:
@@
where $I(x,y,t)$ represents the illuminance at position $(x,y)$ at time $t$. To simplify calculations, assuming consistent global illumination, so $\frac{\partial I}{\partial t}$ is 00, and Eq. 5 simplifies to:
@@
Here, $\nabla I$ represents the spatial gradient of illuminance at the specified position $(x,y)$.
@@
$log\mathcal{I}(x,y,t)-log\mathcal{I}(x,y,t-\Delta t)=pC,$ (7)
where $\Delta t$ is the time interval between consecutive events and $C$ is the contrast threshold of the event camera. This equation indicates that an event triggers when the logarithmic illuminance change at a location exceeds the threshold $C$. Combining Eqs. 6 and 7, it can be observed that
@@
Point position updates rely on consistent appearance feature matching over time. However, velocity variations disrupt the stability of event temporal distribution, causing significant matching errors. To tackle the problem, VMAM is introduced to guide appearance feature matching using kinematic features extracted through MGM. Specifically, to obtain the correlation map $c_{t}^{k}$ at the $k$-th iteration and time $t$, VMAM samples point features from the initial, $t-4$ and $t-2$ time at corresponding coordinates. These features are concatenated and fused with temporally contextual kinematic features via cross-attention. The fused features are then divided into two branches: one captures short-term temporal dependencies via 1D temporal convolution, while the other extracts long-term dependencies through temporal attention. The short-range and long-range features are summed together and correlated with the spatial context features of $x_{t}^{k}$ at time $t$, yielding $c_{t}^{k}$.

### v2

# MATE: Motion-Augmented Temporal Consistency for Event-based Point Tracking

@@ Abstract
Tracking Any Point (TAP) plays a crucial role in motion analysis. Video-based approaches rely on iterative local matching for tracking, but they assume linear motion during the blind time between frames, which leads to point loss under large displacements or nonlinear motion. The high temporal resolution and motion blur-free characteristics of event cameras provide continuous, fine-grained motion information, capturing subtle variations with microsecond precision. This paper presents an event-based framework for tracking any point, which tackles the challenges posed by spatial sparsity and motion sensitivity in events through two tailored modules. Specifically, to resolve ambiguities caused by event sparsity, a motion-guidance module incorporates kinematic vectors into the local matching process. Additionally, a variable motion aware module is integrated to ensure temporally consistent responses that are insensitive to varying velocities, thereby enhancing matching precision. To validate the effectiveness of the approach, two event dataset for tracking any point is constructed by simulation. The method improves the $Survival_{50}$ metric by 17.9% over event-only tracking of any point baseline. Moreover, on standard feature tracking benchmarks, it outperforms all existing methods, even those that combine events and video frames.
@@ 1 Introduction
Figure 2: (a) The spatio-temporal distribution of events generated by a stick rotating uniformly around a pivot. Sampling from different regions along the stick at fixed time intervals reveals spatial sparsity. The event count increases with distance from the pivot, corresponding to higher linear velocity. This indicates that faster motion triggers more events, and vice versa. (b) This paper leverages the temporal continuity of events to mitigate spatial sparsity and exploits subtle motion to guide matching, ultimately improving tracking accuracy.
Recent methods rely on video input, predicting the positions of query points by matching their appearance features with local regions in subsequent frames [9, 10, 16]. However, as these methods assume slow linear motion during the blind time between frames, they face the challenge of objects may undergo large displacements or nonlinear motion, causing query points to exceed the bounds of local regions and resulting in ambiguities in feature matching, see Fig. 1. While some approaches attempt to mitigate this by considering spatial context [39, 7], they still struggle to overcome the lack of motion during blind time.
To cope with the above issue, this paper utilize event cameras to capture the motion during blind time. Event cameras are bio-inspired sensors that respond to pixel-level brightness changes with microsecond temporal resolution, generating sparse and asynchronous event streams [24, 40]. They have the characteristics of high temporal resolution, no motion blur, and low energy consumption. Therefore, parsing motion during blind times using event streams is a feasible solution for tracking any point. Furthermore, capturing motion alone significantly reduces computational overhead compared to traditional frame-based cameras, enabling more efficient methods.
Recent methods [25, 14] have attempted to use event cameras for tracking any point. However, they still rely solely on appearance-based matching for tracking. As shown in Fig. 2(a), the unique spatio-temporal properties of events make existing methods difficult to apply, manifesting in two ways: 1) Event cameras respond only to pixels with brightness changes, resulting in a sparse spatial distribution that introduces ambiguities in appearance-based matching due to spatial discontinuities. 2) The variable speed of scene objects causes fluctuations in event update frequencies, impacting the temporal consistency of event representations. Higher motion speeds lead to increased event update frequencies and densities, and vice versa. Density variations cause inconsistencies in event representations over time, affecting the precision and reliability of temporal matching.
@@
Therefore, this paper proposes a novel framework that utilizes Motion-Augmented Temporal consistency for Event-based point tracking (MATE). It includes a Motion-Guidance Module (MGM) and a Variable Motion Aware (VMA) module. Specifically, MGM leverages the gradient from event time surface to compute kinematic features that construct a dynamic-appearance space, clarifying ambiguities in feature matching and guiding the extraction of temporally consistent appearance features. Additionally, to model the non-stationary states of events, VMA is introduced to employ kinematic cues for adaptive correction and generate temporally consistent responses by combining long-term memory parameters with hierarchical appearance. The method is evaluated across four datasets on two tasks, with experimental results demonstrating its superiority. The main contributions of this paper can be summarized as follows:
  • An event-based framework for tracking any point, named MATE, is presented to monitor surface points on objects by leveraging the temporal continuity of events.
  • This paper reveals the impact of spatial sparsity and motion sensitivity in events on TAP. To address these limitations, a motion-guidance module is proposed to enhance matching process using temporal continuity, while a variable motion aware module models kinematic cues to estimate non-stationary event states, improving performance.
  • Experimental results demonstrate that the proposed approach outperforms current state-of-the-art methods in both tracking of any point and feature tracking tasks across four datasets, with fewer model parameters.
@@ 2.1 Event-based Optical Flow
Figure 3: MATE Overview. At the initial stage, given query points and event data, the encoder extracts feature representations $\mathbf{F}$. At the $k$-th iteration, motion guidance module derives kinematic vectors $\mathbf{V}^{k}$ from the time surface to guide appearance matching in variable motion aware module and enrich the correlation map, which is fed into the transformer to predict the next point location. Variable motion aware module extracts hierarchical features from $\mathbf{F}$, utilizes $\mathbf{V}^{k}$ as guidance to compute the correlation map $\mathbf{C}^{k}$. The transformer takes Point position $\mathbf{X}^{k}$, kinematic vectors $\mathbf{V}^{k}$ and correlation map $\mathbf{C}^{k}$ as input to compute the displacement $\Delta\mathbf{X}$ for the next iteration, and this process repeats until the final trajectory is obtained.
@@ 2.2 Event-based Feature Tracking
Feature tracking aims to predict the trajectories of keypoints. Early approaches can be grouped into two types: one [22, 45] treats feature points as event sets and tracks them using the ICP [5] method, while the other [11] extracts feature blocks from reference frames and matches them by computing brightness increments from events. Moreover, event-by-event trackers [2, 3, 19, 38] exploit the inherent asynchronicity of events. Unfortunately, these methods involve complex parameters that require extensive manual tuning for different event cameras and new environments.
@@ 2.3 Tracking Any Point
Video-based tracking of any point methods [9, 10, 16, 7, 44] have been developed for a few years. These methods model the appearance around the points, using MLPs to capture long-range temporal contextual relationships across frames. During inference, they employ a sliding time window to handle long videos. For example, PIPs [16] frames pixel tracking as a long-range motion estimation problem, updating trajectories through iterative local searches. In contrast, TAP-Net [9] formalizes the problem as tracking any point, overcoming occlusion through global search. However, these methods track points independently, leading to ambiguities in feature matching. Consequently, some studies [39, 7, 21, 42, 41] have been proposed to utilize spatial context to alleviate this issue. In addition to methodological innovations, the PointOdyssey [44] and the TAP-Vid-Davis [9] datasets are collected to advance the field and are used as training data or benchmark datasets for evaluation.
Event-based tracking of any point methods [25, 14] have only been proposed in the past year. FE-Tap [25] performs tracking of any point by fusing video frames and events, while ETAP [14] constructs a backward event flow and optimizes network parameters using contrastive loss.
Although these methods have achieved promising results, they all rely on appearance features for correspondence matching. However, the spatial sparsity of events leads to misalignment when solely modeling based on appearance. Additionally, the temporal inconsistency in event density, caused by variable-speed motion, affects temporal context modeling. In this paper, a motion-guidance module is designed to construct a dynamic-appearance matching space, thereby reducing matching ambiguity. Moreover, a variable motion aware module is employed to generate temporally consistent responses for correlation operations. The method is trained and tested on simulated event modalities derived from the [44] dataset.
@@ 3.1 Setup and Overview
Given a query point $x_{src}\in\mathbb{R}^{2}$ at the initial time, subsequent events are represented by the Time Surface (TS) [29], where each pixel records the timestamp of the most recent event, capturing motion process over a period of time. During the initialization period, the TS representation is fed into a residual network [17] to extract features $\mathbf{F}^{T\times C\times H\times W}=\left\{\mathbf{F}_{0},\mathbf{F}_{1},...,\mathbf{F}_{T}\right\}$, where $T,C,H,W$ represent the number of time steps, channel size, height, and width of feature maps, respectively.
At the iterative stage, let $x_{t}^{k}$ denote the coordinate of the query point at time $t$ after the $k$-th iteration. The MGM estimaties the parameters of the local timestamp plane around $x_{t}^{k}$, yielding the kinetic vector $v_{t}^{k}$, which encodes the local motion tendency. To compute the displacement of coordinate, the VMA module samples local feature patches from $\mathbf{F}$ at the coordinates $\{x_{0}^{k},x_{t-4}^{k},x_{t-2}^{k}\}$, constructing feature pyramids $\{f_{0}^{k},f_{t-4}^{k},f_{t-2}^{k}\}$. These patches are aligned under the guidance of $v_{t}^{k}$ and used to compute the correlation map $c_{t}^{k}$ with the feature at $x_{t}^{k}$. The complete correlation set over all time steps is denoted as $\mathbf{C}^{k}=\{c_{0}^{k},c_{1}^{k},...,c_{t}^{k}\}$. The kinetic vectors $\mathbf{V}^{k}$, correlation maps $\mathbf{C}^{k}$, and point positions $\mathbf{X}^{k}$ are processed by a transformer to predict the displacement $\Delta\mathbf{X}$. The updated coordinates for the next iteration are given by Eq. 1:
$\centering\mathbf{X}^{k+1}=\mathbf{X}^{k}+\Delta\mathbf{X}.\@add@centering$ (1)
@@ 3.2 Motion-Guidance Module
To reduce ambiguities arising from appearance-only matching, a motion-guidance module is designed to leverage the gradient characteristics of the TS to compute kinematic vectors, thereby building a dynamic-appearance matching space and subsequently guiding the extraction of temporally consistent appearance features.
At the first iteration, the query points at all time steps are initialized as $x_{src}$. The MGM samples the TS at $x_{src}$ across all time steps to compute the the initial kinetic vectors $\mathbf{V}^{0}\in\mathbb{R}^{T\times 2}$, which represent the motion trend in the $x$- and $y$-directions at each time step. At the $k$-th iteration, MGM samples the TS at $\mathbf{X}^{k}$ across time to update the kinetic vectors $\mathbf{V}^{k}\in\mathbb{R}^{T\times 2}$. Specifically, the event stream after TS encoding is visualized as a surface in the $xyt$ spacetime domain, representing the active events surface $\Sigma_{e}$ [4]. The spatial gradients of this surface describe the temporal changes relative to spatial variations, establishing a derivative relationship with pixel displacement at corresponding positions, see Eq. 2
@@
For the query point $x_{t}^{k}$, this paper treats the surface formed by neighboring pixels as a surface of active events. Using the coordinates and of neighboring points $(x_{1},y_{1},t_{1})$, $(x_{2},y_{2},t_{2})$, …, a system in Eq. 3 is constructed:
@@
Here, $(a,b,c,d)$ are coefficients for the tangent plane. The spatial gradient of the surface is estimated through plane fitting using SVD, providing motion trend at the query point. However, in the presence of overlapping moving objects, the pixels at boundary intersections can exhibit multiple distinct motion states, disrupting local smoothness and resulting in deviations in the kinematic vectors. To address this, the proposed motion-guidance module employs a multi-layer perceptron to capture the temporal motion relationships, dynamically assigning weights to the kinematic vectors at different time steps, thereby correcting the motion ambiguity at the boundaries.
The corrected kinematic vectors serve two key roles: first, as input to VMA, guiding the generation of speed-insensitive appearance features for correlation; second, as inputs to the transformer, combining with correlation maps and point positions to construct a dynamic-appearance space for iterative point displacement updates.
@@ 3.3 Variable Motion Aware Module
Objects in the scene exhibit variable speeds in three main ways: 1) different objects move at distinct speeds, 2) the speed of a single object changes over time, and 3) different parts of the same object move at different speeds. The variation in speed affects the frequency of event updates. Assuming an object moves at speed $\mathbf{v}=(v_{x},v_{y})$, the illuminance change rate at any point on the edge is given by:
@@
[…]
@@
[…]
@@
$logI(x,y,t)-logI(x,y,t-\Delta t)=pC,$ (6)
where $\Delta t$ is the time interval between consecutive events and $C$ is the contrast threshold of the event camera. This equation indicates that an event triggers when the logarithmic illuminance change at a location exceeds the threshold $C$. Combining Eqs. 5 and 6, it can be observed that
@@
Point position updates rely on consistent appearance feature matching over time. However, velocity variations disrupt the stability of event temporal distribution, causing significant matching errors. To tackle the problem, VMA is introduced to guide appearance feature matching using kinematic vectors extracted through MGM.
Specifically, to obtain the correlation map $c_{t}^{k}$ at the $k$-th iteration and time $t$, VMA samples patch features from $\mathbf{F}_{0}$, $\mathbf{F}_{t-4}$, $\mathbf{F}_{t-2}$ at {$x_{0}^{k}$, $x_{t-4}^{k}$, $x_{t-2}^{k}$} to construct feature pyramids. After interpolation, the feature pyramid maintains consistent spatial dimensions. The features are then flattened into a one-dimensional vector and concatenated along the channel dimension to form $g_{t}^{k}\in\mathbb{R}^{1\times C}$. Repeating this across all time steps yields $\mathbf{G}^{k}=\{g_{0}^{k},g_{1}^{k},...,g_{t}^{k}\}\in\mathbb{R}^{T\times C}$. These features are fused with temporally contextual kinematic vectors $\mathbf{V}^{k}$ via cross-attention, then processed through two complementary branches: one captures short-term dependencies via two 1D temporal convolutions, while the other extracts long-term dependencies through temporal attention. The short-range and long-range features are summed together and correlated with the spatial context features of $\mathbf{X}^{k}$, producing $\mathbf{C}^{k}$. This means that the correlation map $c_{t}^{k}$ is determined by computing the correlation between the features from time steps {0, $t-4$, $t-2$} and $t$. The reason for sampling at these time is that if tracking was successful on one of these offset frames, the feature will reflect the updated appearance of the query point, yielding a more useful correlation map than the one from $\mathbf{F}_{0}$.